OpenStudy (anonymous):

The coordinates of rectangle ABCD are A(2, -7), B(8, -7), C(8, -10), D(2, -10). What are the coordinates of the point of intersection of the diagonals of the rectangle after the rectangle is reflected across the x -axis? (-5, -8.5) (5, 8.5) (5, -8.5) (-5, 8.5)

OpenStudy (anonymous):

find the mid points of A and C \[\left( \frac{ x1+x2 }{ 2 } ,\frac{ y1+y2 }{ 2 }\right)\] then to find the reflection across x-axis change the sign of y coordinate.

OpenStudy (anonymous):

so i would do 2+8/ 2 -7+-10/2 = (5,-8.5) right ? or do i still change the sign ?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

to find the reflection you have to change the sign.